OpenStudy (ineedhelpnow):

the perimeter of a tennis court is 228ft. 7 times the courts length plus 4 times its width is 690ft. find the length and width of court.

OpenStudy (anonymous):

2(l+w)=228 l+w=228/2=114 l+w=114 7l+4w=690 now solve

OpenStudy (anonymous):

{{L -> 78, w -> 36}}